# A Novel Knapsack-based Financial Portfolio Optimization using Quantum Approximate Optimization Algorithm

### v2

@@ 4.1 Overall Architecture
This section explains how our method processes everything from stock selection to finding the optimal solution. Figure 1 shows that our structured workflow employs a quantum computing approach to portfolio optimization. Initiating with historical stock price data over time, Algorithm 1 calculates the expected returns of the portfolio using the Markowitz portfolio optimization model. These expected returns are crucial in constructing the portfolio as a binary knapsack problem, incorporating computed mean returns, asset weights, and the portfolio’s capacity constraints. This problem is optimized using Algorithm 2, where the approach for optimizing the knapsack-based portfolio optimization problem is employed. At this stage, we prepared the feasible oracle with quantum Fourier transform (QFT) [35] for conditional checking of the feasibility of our proposed solution to the knapsack problem before constructing quantum walk mixer (QWM), as discussed in [33] with QAOA (QWM–QAOA). The result is a bit string revealing which stocks are chosen or omitted from the portfolio. For example, a solution of (1,0) in the portfolio optimization for Stocks A and B means that Stock A is included and Stock B is excluded, thus outlining the final investment strategy.
@@
Algorithm 1 Turn portfolio optimization problem into a binary knapsack algorithm
INPUT $start$: start date of stock prices, $end$: end date of stock prices, $stocks$: list of stocks
OUTPUT $problem$: portfolio optimization as a knapsack problem
1:procedure EncodeToKnapsack($start$, $end$, $stocks$)
2:     $prices$ = HistoricalReturn($start$, $end$, $stocks$)
3:     //Calculate the Mean Return
4:     $E(R)$ = ExpectReturn($prices$)
5:     // Set the weights of each stock to 1
[…]
7:         $weights[i]$ = 1
8:     end for
9:     // Define the capacity of the knapsack
10:     $C$ = Length($E(R)$)/2
11:     //Formulate to a knapsack
12:     $problem$ = ($E(R)$, $weights$, $C$)
13:      return $problem$
14:end procedure
Algorithm 1 was initiated by extracting the historical returns of each stock within a specified time frame, ranging from the initiation to the termination dates. This retrieval of historical returns was achieved using the $HistoricalReturn$ function. Subsequently, the algorithm applies the Markowitz model using the $ExpectReturn$ function, to compute the expected returns denoted as $E(R)$ based on the historical return prices. This variable provides valuable insights into the potential profitability of the selected stocks in the portfolio. In addition, it constructs a covariance matrix serving as an indicator of the correlation in movements between pairs of stocks. From lines 5125125\-12512, the calculated returns and covariance are transformed to represent the values, weights, and constraints, which are essential components for formulating the knapsack problem. Herein, we set the values to the expected returns ($E(R)$), and the weights for each stock were uniformly set to 1, indicating an equal weight allocation. The algorithm also defined the constraints or capacity for the knapsack problem, with $C$ representing half the length of the expected return array ($Length(E(R))/2$).
Algorithm 2 Optimize knapsack with the QWM-QAOA algorithm
INPUT $problem$: knapsack problem , $p$: layer of circuit, $m$: Trotter step count for QAOA
OUTPUT $ar$: approximation ratio representing optimal choice selection
1:procedure OptimizeKnapsack($problem$, $l$, $m$)
2:     // Determine the number of choices from the problem weights
3:     $choices$ = Length($problem.weights$)
4:     //Calculate the total knapsack weight
5:     $total\_weights$ = TotalWeight($problem.weights$)
6:     // Setup quantum registers for choices and weights
7:     $c_{reg}$ = QuantumRegister($choices$)
8:     $w_{reg}$ = QuantumRegister($total\_weights$)
9:     // Initialize the ancillary quantum register
10:     $a_{reg}$ = QuantumRegister($3$)
11:     // Define quantum mixing and phase operations
12:     $H_{m}$= QWalkMixer($problem$,$m$, $c_{reg}$, $w_{reg}$, $a_{reg}$)
13:     $H_{c}$ = PhaseHamiltonian($c_{reg}$, $problem$)
14:     // Start with a superposition state
15:     $|\psi_{0}\rangle$ = |$\ket{+}^{\otimes choices}$
16:     for $k=1$ to $p$ do
17:         if $k\leftarrow 1$  then
18:              $\gamma_{k},\beta_{k}$ = Initialize()
19:         else
20:              ($\gamma_{k}$,$\beta_{k}$) = ($\gamma_{opt},\beta_{opt}$)
21:         end if
22:         // Apply the unitary operators to evolve the state
23:         $U_{B}(\beta_{k})$ = $e^{-i\beta_{k}H_{M}}$
24:         $U_{c}(\gamma_{k})$ = $e^{-i\gamma_{k}H_{c}}$
25:         |$\ket{\psi_{k}}=U_{c}(\gamma_{k})U_{B}(\beta_{k})\ket{\psi_{k-1}}$
26:         // Define the QAOA cost function and Optimize it
27:         $cost(\gamma,\beta)$=$\bra{\psi_{k}}H_{c}\ket{\psi_{k}}$
28:         ($\gamma_{opt},\beta_{opt}$)= ShgoOptimizer($cost(\gamma,\beta)$)
29:     end for
30:     // Measure the final state to get probabilities
31:     $probabilities$ = MeasureProbabilities(|$\ket{\psi_{p}}$
32:     $ar$ = GetOptimalSolution($probabilities$)
33:     return $ar$
34:end procedure
After formulating portfolio optimization into a knapsack problem, we optimized it in Algorithm 2. In this stage, we used QWM–QAOA to prepare the mixing Hamiltonian of QAOA. The steps in lines 2102102\-10210 involve setting up quantum registers and other preparations necessary for QAOA using a function called $QuantumRegister$. Line $12$ focuses on the preparation of the Mixing Hamiltonian $H_{m}$ using the QWM-QAOA approach; this is realized via the $QwalkMixer$, which requires parameters, such as a predefined trotter step count $m$, the number of quantum registers of choices, weights, and ancillaries. In line 14, we prepare the phase Hamiltonian $H_{c}$, which is a vital component for constructing QAOA, using $PhaseHamiltonian$. The specifications of each method’s construction are presented in Subsection 4.3.
At each circuit layer $p$, the algorithm computed the expectation value (line $28$), facilitating the determination of a new set of optimal $2p$ parameter $(\gamma_{opt},\beta_{opt})$. These optimal parameters were identified using the SHGO method and implemented as $ShgoOptimizer$ in line $29$. Finally, the algorithm measured probabilities (line $31$) through a measurement process. These probabilities were used to determine the optimal solution using the $OptimalChoice$ function to determine the best selection of items represented as $ar$ in line $32$.
@@ 4.2 Knapsack-based Portfolio Formulation
Our methodology is based on the fundamental principles of mean-variance optimization, focusing on the Markowitz model. This model historically maximizes the anticipated yield of a portfolio while considering a predetermined level of risk, quantified by the variance of portfolio returns. We denoted the expected return $E(R_{i})$ of stock $i$ by using the mean-variance calculation. Next, we transformed the expected return into a knapsack problem. Subsequently, we extended this model by reinterpreting the expected return as a component of a knapsack problem, thereby aligning portfolio optimization with knapsack problem dynamics; this adaptation involves redefining the expected return of each stock as a value optimized under the constraints of total portfolio risk. This mathematical transformation, which effectively converts portfolio optimization into a knapsack problem, is expressed as follows:
@@ 4.3.1 Feasibility Oracle
The feasibility oracle is used as a hypothetical subroutine that instantly determines whether a proposed solution to the knapsack problem violates any constraints [36]. We can explore solutions within a well-defined space of bitstrings by representing our portfolio model as a binary knapsack problem. We defined $K(N)=(0,1)^{N}$ as the set of all possible bitstrings of length $N$ representing potential portfolio choices.
Furthermore, each possible choice of any of the $N$ items is represented by a bitstring $x\in K(N)$. ). Thus, the subset feasible solution was denoted as $F$ for the knapsack problem, and the feasibility function was defined as
@@
Considerably, the feasibility oracle to be unitary $U_{f}$ can be described as:
@@
We begin by illustrating how $U_{1}$ is prepared. Quantum Fourier Transform (QFT) [35] is employed within the unitary operation $U_{1}$ to facilitate the process of the adding of weights in the knapsack problem.
@@
where we use the notaion $e(t)=exp(2\pi it)$, and $k\in\mathbb{Z}$ is the binary representation of $k_{n}-1,\cdot\cdot\cdot,k_{0}$. Furthermore, weights were added to $QFT\ket{K_{w}}$ using controlled operations applying phase rotations. These operations are conditioned on the bits of the bitstring $x$ indicating the selection of items. Each bit in $x$ determines whether the corresponding weight $w_{i}$ should be added to $K_{w}$. The addition in the Fourier space was related to phase rotation and implemented by a sequence of controlled phase gates. The magnitude of the rotation was determined by $wi$ and the position of the bit controlling the operation. The implementation of controlled additions can be described as follows: for each selected item (where the corresponding bit in $x$ is |$\ket{1}$, a controlled phase rotation is applied to $K_{w}$. Additionally, the phase added to each computational basis state |$\ket{k}$ within $K_{w}$ is proportional to $w_{i}$:
@@
Figure 2 shows how $U_{1}$ is implemented using an algorithm based on QFT.
Figure 2: Circuit implementation of $U_{1}$, which encodes the total weights of item selections into the ancillary register $K_{W}$ using quantum Fourier transform-based arithmetic, adding a predefined offset $C_{0}$ to facilitate subsequent feasibility checking.
Following the implementation of the unitary operation $U_{1}$, $U_{2}$ is described, which comprises a quantum circuit that conditionally modifies the state of an ancillary qubit concerning the sum of weights represented in the register $K_{w}$. Referring to Equation 17, the ancillary qubit |$\ket{y}$ underwent a state flip to signal a valid configuration when the total weight $w(x)$, augmented by a constant offset $C_{0}$ is less than the predefined threshold $C+C_{0}+1$. This transformation was executed through several multi-controlled-NOT gates, where each gate was influenced by a qubit distinct from the weight register. The successful transition of the ancillary qubit’s state post $U_{2}$ signified a feasible solution, adhering to the knapsack’s capacity, which effectively segregates the solution space into permissible and impermissible weights.
Figure 3: Circuit implementation of $U_{2}$, executing a conditional check based on the total weight calculated by $U_{1}$ using several multi-controlled NOT gates to flip the ancillary qubit if the weight condition is satisfied.
@@ 4.3.2 Quantum Walk Mixer for Enforcing Constraints
where Ham $(x,x^{\prime})$ represents the Hamming distance between two binary strings $x$ and $x^{\prime}$, which is the count of positions where the corresponding bits differ. The exact implementation of the desired mixing operator $U_{B}(\beta)$ presents challenges due to its Hamiltonian, requiring additional resources and noncommuting elements. To address this, we employ an alternative operator $\tilde{B}$, closely resembling $B$. This operator is constructed from $V_{i}$, and the inverse $V_{i}^{\dagger}$ for encoding feasibility information of neighboring states into auxiliary qubits that enable the controlled state manipulation. Additionally, the operator includes single-qubit $X_{i}$ gates for creating neighboring state exploration. Moreover, $U_{f}$ the feasible oracle is included to determine the feasibility of states by ensuring that only valid states are mixed. Finally, $R_{i}^{X}(2\beta)$ aids in adjusting the amplitudes of states based on feasibility and further refining the mixing process, whereas auxiliary qubits are employed to store feasibility information.
Figure 4: General circuit for a single-qubit quantum walk mixer illustrating the core operations involved in mixing a quantum state with its feasible neighbor, guided by feasibility oracles and single-qubit rotations, as part of a constrained optimization algorithm.
@@
$V_{i}^{\dagger}\ket{x,f(n_{i}(x)),f(x),f_{i}(x)}$ =|$=\ket{n_{i}(x),f(x)\oplus f(n_{i}(x)),f(x)\oplus f(n_{i}(x)),f_{i}(x)\oplus f_{i}(x)}$ (27)
={|$=\begin{cases}\ket{n_{i}(x),0,0,0},&\text{if }f(x)=f(n_{i}(x)),\\ \ket{n_{i}(x),1,1,0},&\text{otherwise}\end{cases}.$
@@
$[6pt]$ $=(B\otimes\mathbb{I}^{\otimes 3})\ket{x,0,0,0},\quad\forall x\in K(N).$
@@
$[7pt]$ $=\left(U_{B}(\beta)\otimes\mathbb{I}^{\otimes 3}\right)\ket{x,0,0,0}$
where $\tilde{B}$ is used to generate $U_{B}$. Moreover, $\tilde{B}$ is not commute and should use three ancilla qubits. Therefore, the approximate implementation of $e^{-i\beta\tilde{B}}$ needs to use the trotter product formula. Meanwhile, using the identity $R_{i}^{X}(2\beta)$, the state representing the rest of the circuit in Figure 4 is given by
@@
$[3pt]$ $=V_{i}^{\dagger}\left(R_{i}^{X}(2\beta)\otimes\mathbb{I}^{\otimes 2}\otimes\ket{1}\bra{1}+\mathbb{I}^{\otimes(N+2)}\otimes\ket{0}\bra{0}\right)V_{i}$
@@
$e^{-\beta\tilde{B}}=$ $\left(\prod_{i=0}^{N-1}V_{j}^{\dagger}\left(R_{i}^{X}\left(\frac{2\beta}{m}\right)\otimes\mathbb{I}^{\otimes 2}\otimes\ket{1}\bra{1}+\mathbb{I}^{\otimes(N+2)}\otimes\ket{0}\bra{0}\right)V_{j}\right)^{m}$ (35)
$[6pt]$ $+O\left(\frac{N\beta^{2}}{m}\right).$

### v3

@@ 4 Knapsack-based Portfolio Optimization
In this section, we explain our overall architecture, knapsack-based portfolio formulation, and QAOA for the knapsack problem.
@@ 4.1 Overall Architecture
This section outlines the complete process from stock data preprocessing to determining the optimal portfolio configuration using a quantum-enhanced approach. As shown in Figure 1, the workflow begins by processing historical stock prices to compute the daily returns using the Markowitz-based portfolio optimization model. Algorithm 1 calculates the expected returns (E(R)), which are crucial inputs for formulating the portfolio selection as a knapsack problem. This step maps the expected returns, stock weights, and capacity constraints into a knapsack framework to define the feasible portfolio configurations. In the next phase, Algorithm 2 solves the knapsack-based portfolio optimization problem. Here, we first construct the feasible oracle using Quantum Fourier Transform (QFT) [40] to ensure that only valid solutions are considered for further optimization. Once a feasible solution is identified, the Quantum Walk Mixer (QWM) is applied within the Quantum Approximate Optimization Algorithm (QAOA), following the approach of  [38]. The combination of QWM and QAOA (QWM-QAOA) efficiently explores the solution space, minimizing computational complexity while improving the accuracy of the optimization. The final output is a binary string, where each bit represents whether a stock is included or excluded from the portfolio. For instance, a result of (1,0) for two stocks indicates that Stock A is selected, while Stock B is excluded, thus guiding the final portfolio selection and investment strategy.
@@ 4.2 Knapsack-based Portfolio Formulation
Our methodology is based on the fundamental principles of mean-variance optimization, focusing on the Markowitz model. This model historically maximizes the anticipated yield of a portfolio while considering a predetermined level of risk, quantified by the variance of portfolio returns. We denoted the expected return $E(R_{i})$ of stock $i$ by using the mean return calculation. Next, we transformed the expected return into a knapsack problem. Subsequently, we extended this model by reinterpreting the expected return as a component of a knapsack problem, thereby aligning portfolio optimization with knapsack problem dynamics; this adaptation involves redefining the expected return of each stock as a value optimized under the constraints of total portfolio risk. This mathematical transformation effectively converts portfolio optimization into a knapsack problem:
@@
This process can be summarized as Algorithm 1. It begins by extracting the historical returns of each stock within a specified time range, from the start to the end date, using the $HistoricalReturn$ function. From line $5-12$, these historical returns are then passed into the Markowitz model through the $ExpectReturn$ function, which calculates the expected returns, denoted as $E(R)$. These expected returns provide key insights into the potential profitability of the selected stocks in the portfolio problem. Specifically, the expected returns ($E(R)$) are assigned as the values in the knapsack problem, and the weights for each stock are uniformly set to 1, representing equal weighting for all assets. Afterward, the algorithm also defines the capacity constraint ($C$) of the knapsack problem, which is set to half the number of stocks, calculated as $Length(E(R))/2$. This framework effectively transforms the portfolio optimization problem into a knapsack problem formulation. This transformation allows for a more structured handling of assets and makes it possible to apply powerful combinatorial optimization techniques.
Algorithm 1 Formulate Portfolio Optimization as a Binary Knapsack Problem
INPUT: $start$: start date of stock prices, $end$: end date of stock prices, $stocks$: list of stocks
OUTPUT: $problem$: portfolio optimization modeled as a knapsack problem
1:procedure EncodeToKnapsack($start$, $end$, $stocks$)
2:     // Extracte historical return
3:     $prices\leftarrow$ HistoricalReturn($start$, $end$, $stocks$)
4:     // Calculate the expected return
5:     $E(R)\leftarrow$ ExpectReturn($prices$)
6:     // Set a weight of 1 to each stock
7:     for $i\leftarrow 0$ to Length($E(R)$do
8:         $weights[i]\leftarrow 1$
9:     end for
10:     // Set the capacity of knapsack
11:     $C$ = Length($E(R)$) / 2
12:     // Formulate the problem
13:     $problem\leftarrow(E(R),weights,C)$
14:     return $problem$
15:end procedure
@@ 4.3 QAOA for the knapsack problem
After formulating the portfolio optimization as a knapsack problem, we optimized it using Algorithm 2. We employed the QWM-QAOA to prepare the mixing Hamiltonian for the QAOA process. In lines $2-10$, we set up the quantum registers and other necessary preparations for QAOA using the $QuantumRegister$ function. Line $12$ focuses on constructing the Mixing Hamiltonian $H_{m}$ using the QWM-QAOA approach, which is achieved via the $QWalkMixer$ function. This step requires parameters such as a predefined trotter step $m$, the number of quantum registers, weights, and ancillaries. Next, the algorithm prepares the phase Hamiltonian $H_{c}$ using $PhaseHamiltonian$ (line $13$), which plays a crucial role in the construction of the QAOA. More details on these concepts will be provided in the following sections. At each circuit layer $p$, the algorithm computed the expectation value (line $28$), facilitating the determination of a new set of optimal $2p$ parameter $(\gamma_{opt},\beta_{opt})$. These optimal parameters were identified using the SHGO optimization method and implemented as $ShgoOptimizer$ line $29$. Finally, the algorithm measured probabilities (line $31$) through a measurement process. These probabilities were used to determine the optimal solution using the $OptimalChoice$ function to choose the best selection of items represented as $ar$ as the optimal choice of approximation ratio.
Algorithm 2 Optimize Knapsack Problem Using the QWM-QAOA Algorithm
INPUT $problem$: knapsack problem, $p$: circuit layer, $m$: trotter step count for QAOA
OUTPUT $ar$: approximation ratio as optimal choice
1:procedure OptimizeKnapsack($problem$, $l$, $m$)
2:     // Determine problem choices
3:     $choices$ = Length($problem.weights$)
4:     // Calculate the total weight
5:     $total\_weights$ = TotalWeight($problem.weights$)
6:     // Setup quantum registers
7:     $c_{reg}$ = QuantumRegister($choices$)
8:     $w_{reg}$ = QuantumRegister($total\_weights$)
9:     // Initialize quantum register
10:     $a_{reg}$ = QuantumRegister($3$)
11:     // Define quantum mixing
12:     $H_{m}$= QWalkMixer($problem$,$m$, $c_{reg}$, $w_{reg}$, $a_{reg}$)
13:     // Define phase operations
14:     $H_{c}$ = PhaseHamiltonian($c_{reg}$, $problem$)
15:     // Start with superposition state
16:     $|\psi_{0}\rangle$ = |$\ket{+}^{\otimes choices}$
17:     for $k=1$ to $p$ do
18:         if $k\leftarrow 1$  then
19:              $\gamma_{k},\beta_{k}$ = Initialize()
20:         else
21:              ($\gamma_{k}$,$\beta_{k}$) = ($\gamma_{opt},\beta_{opt}$)
22:         end if
23:         // Apply the unitary operators
24:         $U_{B}(\beta_{k})$ = $e^{-i\beta_{k}H_{M}}$
25:         $U_{c}(\gamma_{k})$ = $e^{-i\gamma_{k}H_{c}}$
26:         |$\ket{\psi_{k}}=U_{c}(\gamma_{k})U_{B}(\beta_{k})\ket{\psi_{k-1}}$
27:         // Define the QAOA cost function
28:         $cost(\gamma,\beta)$=$\bra{\psi_{k}}H_{c}\ket{\psi_{k}}$
29:         // Optimize the cost function
30:         ($\gamma_{opt},\beta_{opt}$)= ShgoOptimizer($cost(\gamma,\beta)$)
31:     end for
32:     // Measure the final state
33:     $probabilities$ = MeasureProbabilities(|$\ket{\psi_{p}}$
34:     // Obtain the optimal probability
35:     $ar$ = GetOptimalSolution($probabilities$)
36:     return $ar$
37:end procedure
@@ 4.3.1 Feasibility Oracle
The feasibility oracle is used as a hypothetical subroutine that instantly determines whether a proposed solution to the knapsack problem violates any constraints [41]. We can explore solutions within a well-defined space of bitstrings by representing our portfolio model as a binary knapsack problem. We defined $K(N)=(0,1)^{N}$ as the set of all possible bitstrings of length $N$ representing potential portfolio choices. Furthermore, each possible choice of any of the $N$ items is represented by a bitstring $x\in K(N)$. ). Thus, the subset feasible solution was denoted as $F$ for the knapsack problem, and the feasibility function was defined as
@@
Considerably, the feasibility oracle to be unitary $U_{f}$ as:
@@
We begin by illustrating how $U_{1}$ is prepared. Quantum Fourier Transform (QFT) [40] is employed within the unitary operation $U_{1}$ to facilitate the process of adding weights in the knapsack problem.
@@
where we use the notaion $e(t)=exp(2\pi it)$, and $k\in\mathbb{Z}$ is the binary representation of $k_{n}-1,\cdot\cdot\cdot,k_{0}$. Furthermore, weights were added to $QFT\ket{K_{w}}$ using controlled operations applying phase rotations. These operations are conditioned on the bits of the bitstring $x$ indicating the selection of items. Each bit in $x$ determines whether the corresponding weight $w_{i}$ should be added to $K_{w}$. The addition in the Fourier space was related to phase rotation and implemented by a sequence of controlled phase gates. The magnitude of the rotation was determined by $w_{i}$ and the position of the bit controlling the operation. The implementation of controlled additions can be described as follows: for each selected item (where the corresponding bit in $x$ is |$\ket{1}$, a controlled phase rotation is applied to $K_{w}$. The phase added to each computational basis state |$\ket{k}$ within $K_{w}$ is proportional to $w_{i}$:
@@
Figure 2 shows how $U_{1}$ is implemented using an algorithm based on QFT. Following the implementation of the unitary operation $U_{1}$, $U_{2}$ is described in Figure 3, which comprises a quantum circuit that conditionally modifies the state of an ancillary qubit concerning the sum of weights represented in the register $K_{w}$. Referring to Equation 16, the ancillary qubit |$\ket{y}$ underwent a state flip to signal a valid configuration when the total weight $w(x)$, augmented by a constant offset $C_{0}$ is less than the predefined threshold $C+C_{0}+1$. This transformation was executed through several multi-controlled-NOT gates, where each gate was influenced by a qubit distinct from the weight register. The successful transition of the ancillary qubit’s state post $U_{2}$ signified a feasible solution, adhering to the knapsack’s capacity, which effectively segregates the solution space into permissible and impermissible weights.
Figure 2: Circuit implementation of $U_{1}$. It encodes the total weights of item selections into the ancillary register $K_{W}$ using quantum Fourier transform-based arithmetic. A predefined offset $C_{0}$ is added to the register to facilitate subsequent feasibility checking.
Figure 3: Circuit implementation of $U_{2}$. It executes a conditional check based on the total weight calculated by $U_{1}$ using several multi-controlled NOT gates to flip the ancillary qubit if the weight condition is satisfied.
@@ 4.3.2 Quantum Walk Mixer for Enforcing Constraints
where Ham $(x,x^{\prime})$ represents the Hamming distance between two binary strings $x$ and $x^{\prime}$, which is the count of positions where the corresponding bits differ. The exact implementation of the desired mixing operator $U_{B}(\beta)$ presents challenges due to its Hamiltonian, requiring additional resources and non-commuting elements. To address this, we employ an alternative operator $\tilde{B}$, closely resembling $B$. This operator is constructed from $V_{i}$, and the inverse $V_{i}^{\dagger}$ for encoding feasibility information of neighboring states into auxiliary qubits that enable the controlled state manipulation. Additionally, the operator includes single-qubit $X_{i}$ gates for creating neighboring state exploration. Moreover, $U_{f}$ the feasible oracle is included to determine the feasibility of states by ensuring that only valid states are mixed. Finally, $R_{i}^{X}(2\beta)$ aids in adjusting the amplitudes of states based on feasibility and further refining the mixing process, whereas auxiliary qubits are employed to store feasibility information.
Figure 4: General circuit for a single-qubit quantum walk mixer. It illustrates the core operations involved in mixing a quantum state with its feasible neighbor, guided by feasibility oracles and single-qubit rotations, as part of a constrained optimization algorithm.
@@
$V_{i}^{\dagger}\ket{x,f(n_{i}(x)),f(x),f_{i}(x)}$ (26)
=|$\quad=\ket{n_{i}(x),f(x)\oplus f(n_{i}(x)),f(x)\oplus f(n_{i}(x)),f_{i}(x)\oplus f_{i}(x)}$
={|$\quad=\begin{cases}\ket{n_{i}(x),0,0,0},&\text{if }f(x)=f(n_{i}(x)),\\ \ket{n_{i}(x),1,1,0},&\text{otherwise}\end{cases}$
@@
$=(B\otimes\mathbb{I}^{\otimes 3})\ket{x,0,0,0},\quad\forall x\in K(N).$
@@
$=\left(U_{B}(\beta)\otimes\mathbb{I}^{\otimes 3}\right)\ket{x,0,0,0}$
where $\tilde{B}$ is used to generate $U_{B}$. Moreover, $\tilde{B}$ is not commute and should use three ancilla qubits. Therefore, the approximate implementation of $e^{-i\beta\tilde{B}}$ needs to use the trotter product formula. Meanwhile, using the identity $R_{i}^{X}(2\beta)$, the state representing the rest of the circuit in Figure 4.
@@
$=V_{i}^{\dagger}\left(R_{i}^{X}(2\beta)\otimes\mathbb{I}^{\otimes 2}\otimes\ket{1}\bra{1}+\mathbb{I}^{\otimes(N+2)}\otimes\ket{0}\bra{0}\right)V_{i}$
@@
$e^{-\beta\tilde{B}}=$ $|\displaystyle\left(\prod_{i=0}^{N-1}V_{j}^{\dagger}\left(R_{i}^{X}\left(\frac{2\beta}{m}\right)\otimes\mathbb{I}^{\otimes 2}\otimes\ket{1}\bra{1}\right.\right.$ | (34)
$\left.\left.\quad+\mathbb{I}^{\otimes(N+2)}\otimes\ket{0}\bra{0}\right)V_{j}\right)^{m}$
$\quad+O\left(\frac{N\beta^{2}}{m}\right)$